A sled is pulled from rest by a net force of 2.8 N [forward]. The sled moves 7.4 m, reaching a velocity of 3.9 m/s [forward]. Determine the mass of the sled. (determine the mass by using one of the kinematic equations).


1
Expert's answer
2022-05-02T13:36:14-0400

The acceleration of the sled

"a=\frac{v^2-v_0^2}{2d}=\frac{3.9^2}{2*7.4}=1.03\:\rm m/s^2"

The mass of the sled

"m=\frac{F}{a}=\frac{2.8}{1.03}=2.7\:\rm kg"